OpenStudy (anonymous):

Solve the equation for x: 4000=1000e^2x

OpenStudy (anonymous):

4000=1000e^2x e^2x=4 take log both sides log(e^2x)=log4 2x*loge=log4 2x=log4 {bcoz loge=1} 2x=.6020 x=.6020/2=.3010 according to me

OpenStudy (anonymous):

these are my options.... x=ln4-2 x=ln4000/2(ln1000) x=1n4/2 x=4/e^2

OpenStudy (anonymous):

x=log4/2 is the correct one